If a wet sheet in a dryer losses its moisture at a rate proportional to its moisture content, and losses half of its moisture during the first 10 minutes, when it will be practically dry? Say when will it have lost 99% of its moisture.


1
Expert's answer
2021-05-04T13:21:56-0400

"\dfrac{dV}{dt}∝ V"

Where V is Volume of moisture in towel


"\dfrac{dV}{dt }= kV"

Where k is proportionality constant


On Integrating,

"\ln V = kt +C"


At t = 0s, V=Vi

"C =\ln V_i"


At t=600s, V= 0.5Vi

"\ln(0.5V_i) =600k+ \ln V_i\\



\ln(0.5V_i) - \ln V_i = 600k\\"


"\dfrac{\ln(\frac{V_i}{2Vi})}{600 }= k"


"k = \dfrac{\ln(\frac{1}{2})}{600}"


"k = \dfrac{\ln(1) - \ln(2)}{600}"


"k= \dfrac{0 -ln(2)}{600 }= \dfrac{\ln(2)}{600}=\dfrac{ -0.693}{600}"


"k = -1.1552 \u00d7 10^{{-3}}"


According to the question,

"V= V_i - \dfrac{99V_i}{100} = \dfrac{Vi}{100}"


"\ln\dfrac{V_i}{100}= -1.1552×10^{-3} t +\ lnV_i"


"\dfrac{\ln V_i - ln\frac{V_i}{100}}{1.1552×10^{-3}}= t"


"t = \dfrac{\ln(100)}{1.1552×10^{-3}}=\dfrac{4.6051}{1.1552×10^{-3}}"


"t = 3986 s= 66.43\ min = 1.107\ hour"



Note: You can also use mass or moles instead of volume in the calculations because it will still cancel out.